# Learning Minimal NAP Specifications for Neural Network Verification

### v1

@@ 1. Introduction
While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023). For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5. Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
To this end,  Geng et al. (2023) proposes using neural activation patterns (NAPs) as specifications. Neural activation patterns can be viewed as an abstraction of each neuron’s state. For ReLU activation, it can be categorized into an activation state and a deactivation state. Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class. Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1. This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
@@ 2.4. NAP Specifications for Robustness Verification
Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$. Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$. Geng et al. (2023) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications. And finding qualified NAP specifications implies the underlying robustness problem is verified. We formally frame these requirements into the following properties:
@@
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications. Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou. In this paper, we will not dive into NAP encodings. Interested readers can refer to Geng et al. (2023) for more details.
@@ 3.1. Problem Formulation
  $P\vdash\Delta_{1}\Downarrow P^{\prime}$         $P^{\prime}\vdash\Delta_{2}\Downarrow P^{\prime\prime}$              $P\vdash\Delta_{1};\Delta_{2}\Downarrow P^{\prime\prime}$
                           $P_{i,l}=\mathbf{*}$ Refine              $P\vdash\widetilde{\Delta}(N_{i,l})\Downarrow\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$                                 $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ Coarsen                $P\vdash\dot{\Delta}(N_{i,l})\Downarrow\dot{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$
@@ 2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
             return $None$ ;   /* Return None if even the most refined NAP fails verification */
4
@@
                   $P\leftarrow\dot{\mathcal{A}}(N)$   /* Refinement starts from the coarsest NAP $\dot{P}$/
                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$   /* Returns all combinations of size k */
@@
                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$  /* Refine neurons in the chosen combination */
9
@@
                               return $P$   /* Found the minimal NAP */
11
12
@@
             return $None$ ;   /* Return None if the most refined NAP fails verification */
@@
                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;   /* Try to abstract this neuron using $\dot{\Delta}$/
@@
                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;   /* Refine the neuron back if the verification fails */
@@ 4.1. Adversarial-guided Prune
However, when the two NAPs disagree on multiple neurons, things become a little bit different. Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$. We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$. In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification. In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 4(b), 4(c), 4(d). So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$). Formerly, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of mandatory neurons $M$ if satisfies the following condition:
@@
1 Function Adversarial_Prune($N$, $X$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $x_{j}$ in $X$ do
3             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$ for $N_{i,l}$ in $N$ do
@@
                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Create mandatory neurons for $x^{\prime}_{j}$/
5
6
@@ 4.2. Gradient-guided Search
1 Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $N_{i,l}$ in $N$ do
@@ 4                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$ if $\hat{z}_{j}>0$  then
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{0}$, can’t coarsen */
6
@@
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{1}$, can’t coarsen */
9
10
11
12
@@ 5.2. The Statistical Coarsen Approach
1 Function Sample_NAPs($unvisited$, $\theta$)
       $P\leftarrow\dot{\mathcal{A}}(N)$   /* Use the coarsest NAP as a blank template */
@@ 1 Function Sample_NAPs($unvisited$, $\theta$)
3             $rand\leftarrow\text{random}(0,1)$ if $rand\leq\theta$ then
                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$   /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
4
5
@@ 3             $rand\leftarrow\text{random}(0,1)$ if $rand\leq\theta$ then
7Function Sample_Refine($visited$, $\theta$, $k$, $s$)
@@
             $unvisited\leftarrow N\setminus visited$ $ctr=dict()$   /* Create a counter for each neuron in $unvisited$/
9             for _normal-_\__ in $range(k)$ do
                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$   /* Sample $k$ NAPs */
@@
                               $ctr[N_{i,l}]$ += 1   /* Count successful NAPs for each neuron */
12
13
14
            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$   /* Pick potential mandatory neuron */
@@ 15             $visited\leftarrow visited\cup\{N_{mad}\}$
       return $None$   /* Return None if the most refined NAP fails verification */
18
@@ 21             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
             $visited\leftarrow\emptyset$   /* Start from the coarsest NAP */
23
@@ 24      $Sample\_Refine(visited,\theta,k,s)$
1 Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$ if $\mathcal{V}(P)==0$ then
             return $None$   /* Return None if the most refined NAP fails verification */
3
@@ 2       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$ if $\mathcal{V}(P)==0$ then
6                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$ if $\mathcal{V}(P)==1$  then
7                         $found\_neurons\leftarrow\emptyset$ for $N_{i,l}$ in $N$ do
@@ 9                                     $found\_neurons\leftarrow found\_neurons\cup\{N_{i,l}\}$
10
                        $mand\_neurons\leftarrow found\_neurons$   /* Reduce search space */
11
@@
                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$   /* Sample a new NAP */
13
14
            return $P$   /* Return the minimal NAP of size $s$/
15
16
@@ 7. Evaluation
To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset. We specifically focus on classes 0, 1, and 4, as the most refined NAPs for other labels either encounter timeouts or fail the verification. We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$. The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023).
@@ 8.1. Neural Network Verification
Neural network verification has attracted much attention due to the increasingly widespread applications of neural networks in safety-critical systems. it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017) makes it a challenging task. Thus most of the existing work on neural network verification focuses on designing scalable verification algorithms. For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verifying small-scale neural networks with fewer than 100 neurons, state-of-the-art approaches (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks. It is worth mentioning that most existing work adopts local neighborhood specifications to verify the robustness properties of neural networks (Shapira et al., 2023). Despite being a reliable measure, using specifications that define local neighborhood of reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data. Yet evidently, being able to verify the neural network on test data is an important task in real-world applications. The development of NAP specification (Geng et al., 2023) sheds light on this challenge. While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness. In this paper, we challenge this assumption by demonstrating that only a small fraction of neurons are crucially related to robustness. This not only allow us study robustness from the neurons’ perspective but also significantly pushes the known verifiable boundaries.

### v2

@@ 1. Introduction
While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023a). For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5. Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
To this end,  Geng et al. (2023a) proposes using neural activation patterns (NAPs) as specifications. Neural activation patterns can be viewed as an abstraction of each neuron’s state. For ReLU activation, it can be categorized into an activation state and a deactivation state. Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class. Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1. This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
@@ 2.4. NAP Specifications for Robustness Verification
Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$. Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$. Geng et al. (2023a) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications. And finding qualified NAP specifications implies the underlying robustness problem is verified. We formally frame these requirements into the following properties:
@@
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications. Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou. In this paper, we will not dive into NAP encodings. Interested readers can refer to Geng et al. (2023a) for more details.
@@ 3.1. Problem Formulation
[…]
@@ 4       $P\leftarrow\widetilde{\mathcal{A}}(N)$
             return $None$ ;
              /* Return None if even the most refined NAP fails verification */
6
@@
                   $P\leftarrow\dot{\mathcal{A}}(N)$
                    /* Refinement starts from the coarsest NAP $\dot{P}$/
                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$
                    /* Returns all combinations of size k */
@@
                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$
                                /* Refine neurons in the chosen combination */
11
@@
                               return $P$
                                /* Found the minimal NAP */
@@ 2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
             return $None$ ;
              /* Return None if the most refined NAP fails verification */
@@
                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;
                    /* Try to abstract this neuron using $\dot{\Delta}$/
@@
                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;
                          /* Refine the neuron back if the verification fails */
@@ 4.1. Adversarial-guided Prune
However, when the two NAPs disagree on multiple neurons, things become a little bit different. Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$. We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$. In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification. In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 5(b), 5(c), 5(d). So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$). Formerly, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of mandatory neurons $M$ if satisfies the following condition:
@@
1
2Function Adversarial_Prune($N$, $X$)
3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
4       for $x_{j}$ in $X$ do
5             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$
6             for $N_{i,l}$ in $N$ do
@@ 5             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$
                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$
                          /* Create mandatory neurons for $x^{\prime}_{j}$/
8
9
@@ 4.2. Gradient-guided Search
1
2Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
4       for $N_{i,l}$ in $N$ do
@@ 6                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
                                /* Neuron is in state $\mathbf{0}$, can’t coarsen */
9
10
@@
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
                                /* Neuron is in state $\mathbf{1}$, can’t coarsen */
13
14
15
16
@@ 5.2. The Statistical Coarsen Approach
1
2Function Sample_NAPs($unvisited$, $\theta$)
       $P\leftarrow\dot{\mathcal{A}}(N)$
        /* Use the coarsest NAP as a blank template */
@@
4             $rand\leftarrow\text{random}(0,1)$
5             if $rand\leq\theta$ then
                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$
                    /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
6
7
@@ 4             $rand\leftarrow\text{random}(0,1)$
9
10
11Function Sample_Refine($visited$, $\theta$, $k$, $s$)
12
@@
14             $unvisited\leftarrow N\setminus visited$
             $ctr=dict()$
              /* Create a counter for each neuron in $unvisited$/
15             for __\__ in $range(k)$ do
                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$
                    /* Sample $k$ NAPs */
@@ 14             $unvisited\leftarrow N\setminus visited$
                               $ctr[N_{i,l}]$ += 1
                                /* Count successful NAPs for each neuron */
18
19
20
            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$
              /* Pick potential mandatory neuron */
@@ 21             $visited\leftarrow visited\cup\{N_{mad}\}$
27 if $\mathcal{V}(P)==0$ then
       return $None$
        /* Return None if the most refined NAP fails verification */
28
@@ 31             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
             $visited\leftarrow\emptyset$
              /* Start from the coarsest NAP */
33
@@ 34      $Sample\_Refine(visited,\theta,k,s)$
1
2Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
3       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$
4       if $\mathcal{V}(P)==0$ then
             return $None$
              /* Return None if the most refined NAP fails verification */
5
@@ 3       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$
8                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$
9                   if $\mathcal{V}(P)==1$  then
10                         $found\_neurons\leftarrow\emptyset$
11                         for $N_{i,l}$ in $N$ do
@@ 13                                     $found\_neurons\leftarrow found\_neurons\cup\{N_{i,l}\}$
14
                        $mand\_neurons\leftarrow found\_neurons$
                          /* Reduce search space */
15
@@
                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$
                          /* Sample a new NAP */
17
18
            return $P$
              /* Return the minimal NAP of size $s$/
19
20
21
@@ 7. Evaluation
To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset. We specifically focus on classes 0, 1, and 4, as the most refined NAPs for other labels either encounter timeouts or fail the verification. We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$. The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023a).
@@ 8.1. Neural Network Verification
Neural network verification has attracted much attention due to the increasingly widespread applications of neural networks in safety-critical systems. it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017) makes it a challenging task. Thus most of the existing work on neural network verification focuses on designing scalable verification algorithms. For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verifying small-scale neural networks with fewer than 100 neurons, state-of-the-art approaches (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks. It is worth mentioning that most existing work adopts local neighborhood specifications to verify the robustness properties of neural networks (Shapira et al., 2023). Despite being a reliable measure, using specifications that define local neighborhood of reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data. Yet evidently, being able to verify the neural network on test data is an important task in real-world applications. The development of NAP specification (Geng et al., 2023a) sheds light on this challenge. While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness. In this paper, we challenge this assumption by demonstrating that only a small fraction of neurons are crucially related to robustness. This not only allow us study robustness from the neurons’ perspective but also significantly pushes the known verifiable boundaries.